# Turbulence and far-from-equilibrium equation of state of Bogoliubov waves in Bose-Einstein Condensates

### v2

@@ Abstract
Bogoliubov waves are fundamental excitations of Bose-Einstein Condensates (BECs). They emerge from a perturbed ground state and interact nonlinearly, triggering turbulent cascades. Here, we study turbulent BECs theoretically and numerically using the 3D Gross-Pitaevskii model and its wave-kinetic equations. We derive a new Kolmogorov-like stationary spectrum for short Bogoliubov waves and find a complete analytical expression for the spectrum in the long acoustic regime. We then use our predictions to explain the BEC equation of state reported in [Dora et al. Nature 620,521 (2023)], and to suggest new experimental settings.
@@
Turbulence, in the broad sense, is perhaps one of the most striking manifestations of far-from-equilibrium physics in Nature. In classical fluids, it emerges when the forcing scale, at which energy is injected, is well-separated from the one at which it is dissipated by molecular viscosity. Energy is then carried along scales through a turbulent cascade driven by hydrodynamic interactions. An analogous process occurs in nonlinear wave systems, where cascades are driven by the interaction of waves having different wavelengths. When the wave nonlinearity is weak, the Wave Turbulence Theory (WTT) provides a comprehensive theoretical framework describing far-from-equilibrium steady-states Nazarenko (2011). The theory has been successful in describing, for instance, the dynamics of gravitational waves Galtier and Nazarenko (2017), Kelvin waves on superfluid vortices L’vov and Nazarenko (2010), internal gravity waves in the ocean Caillol and Zeitlin (2000); Galtier (2003), Langmuir waves in plasmas Zakharov (1972), and matter waves in turbulent Bose-Einstein Condensates (BECs) Dyachenko et al. (1992); Zhu et al. (2023a).
One of the most interesting and important regimes in BEC turbulence is realized when weak waves propagate on a background of a coherent ground state, the condensate. These are the Bogoliubov waves, whose frequency $\omega$ in the linear approximation is given by the following dispersion relation,
@@
where $k=|{\bf k}|$ is the modulus of the wave vector ${\bf k}$, $c_{s}$ is the speed of sound, and $\xi$ is the healing length of the BEC at which dispersive effects become important. In the absence of vortices, and with a strong condensate, the leading nonlinearity is quadratic, and wave interactions involve only three waves. When the condensate is negligible, the interaction term is cubic and the relevant nonlinear processes are 4-wave interactions. Over the last few years, outstanding progress has been achieved in realizing turbulent BECs experimentally, and they have become an excellent prototype for studying far-from-equilibrium physics.
@@
Turbulent states are driven by external forcing and dissipation, which results in an energy flux $P_{0}$ through scales (which equals the energy injection and dissipation rates). Such fluxes thus play a role equivalent to the one of temperature in equilibrium thermal states. It is then natural to ask if there is a universal far-from-equilibrium equation of state (EoS) relating energy flux and internal observables, such as the energy or the wave-action spectrum. The clean and well-controlled turbulent BEC experiments available today Dogra et al. (2023), together with recent numerical simulations of the Gross-Pitaevskii equation Martirosyan et al. (2024), have led to revealing the existence of such EoS. They have shown that the amplitude of the wave-action spectrum scales with the flux as $P_{0}^{1/3}$, followed by a $P_{0}^{2/3}$-scaling for large flux. Whereas the $P_{0}^{1/3}$-scaling can be understood within the WTT invoking 4-wave interactions, the second scaling-law lacks explanation. Also, data uncertainties do not uniquely fix the exponents.
In this Letter, we will revise old and obtain new theoretical predictions for the BEC turbulence dominated by interacting Bogoliubov waves. We test these predictions by numerical simulations and use our results to provide a plausible theoretical explanation of the second scaling law of the EoS observed in BEC turbulence. In particular, we argue that it is consistent with scaling $P_{0}^{1/2}$ arising from interacting short-wave Bogoliubov modes. We re-analyze the experimental data of Dogra et al. (2023), and show their agreement with our prediction.
@@
The master model for describing BECs is the Gross-Pitaevskii equation (GPE) for the complex scalar wave function $\Psi({\bf x},t)$, where ${\bf x}$ is the position in 3D physical space and $t$ is time. Defining $\Psi$ such that $|\Psi|^{2}$ is the mass density, the GPE written in terms of $\xi$, $c_{s}$, and the unperturbed (uniform) condensate mass density $\rho_{0}$, reads
$i\frac{\partial\Psi}{\partial t}=\frac{c_{s}}{\sqrt{2}\xi}\Big{[}-\xi^{2}\nabla^{2}+\frac{|\Psi|^{2}}{\rho_{0}}-1\Big{]}\Psi+i\,F({\bm{x}},t)+i\frac{\nabla^{2\alpha}\Psi}{k_{\rm R}^{2\alpha}}.$ (2)
We have also included a large-scale forcing $F$ and a hyper-viscous dissipation term acting on small scales, both necessary for describing turbulent states. The healing length and the speed of sound depend both on the physical properties of the BEC and $\rho_{0}$.
@@
The GPE can be mapped to an effective compressible irrotational fluid flow via the so-called Madelung transformation, $\Psi({\bf x},t)=\sqrt{\rho({\bf x},t)}\exp[i\phi({\bf x},t)/\sqrt{2}c_{s}\xi]$ with $\rho({\bf x},t)$ and $\phi({\bf x},t)$ being the fluid mass density and the velocity potential, respectively. The Bogoliubov dispersion relation (1) is obtained by linearizing the GPE (2), in the absence of forcing and dissipation terms, considering perturbations over the ground state, $\tilde{\Psi}=\Psi-\sqrt{\rho_{0}}$ and diagonalizing them in Fourier space. For weakly nonlinear waves, the leading order interaction occurs between the modes satisfying three-wave resonant conditions,
@@
In this limit, the WTT provides a wave-kinetic equation (WKE) describing the evolution of the wave-action spectrum $n_{{\bf k}}=n({{\bf k}},t)$. The general WKE for weakly nonlinear waves driven by three-wave resonant interactions is Zakharov et al. (1992); Nazarenko (2011)
$\frac{\partial n_{{\bf k}}}{\partial t}=\mathrm{St}_{{\bf k}},$ (4)
@@
$\begin{split}&\mathrm{St}_{{\bf k}}=2\pi\int\left(\mathcal{R}^{{\bf k}}_{\bf{12}}-\mathcal{R}^{\bm{1}}_{\bm{k2}}-\mathcal{R}^{\bm{2}}_{\bm{k1}}\right)\mathrm{d}{\bf k}_{1}\mathrm{d}{\bf k}_{2}\,,\\ &\mathcal{R}^{{\bf k}}_{\bm{12}}=\delta^{{\bf k}}_{\bm{12}}\delta(\omega^{k}_{{12}})|V^{k}_{{12}}|^{2}\left[n_{{{\bf k}}_{1}}n_{{{\bf k}}_{2}}-n_{{{\bf k}}}n_{{{\bf k}}_{1}}-n_{{{\bf k}}}n_{{{\bf k}}_{2}}\right]\,,\\ &\delta^{{\bf k}}_{\bf{12}}=\delta({{\bf k}}-{\bf k}_{1}-{\bf k}_{2})\,,\\ &\delta(\omega^{k}_{{12}})=\delta(\omega_{k}-\omega_{{k}_{1}}-\omega_{{k}_{2}})\,,\end{split}$ (5)
where $V^{k}_{{12}}\equiv V({k},{k_{1}},{k_{2}})$ is the three-wave interaction amplitude depending on the particular type of waves. For Bogoliubov waves, they are given by
$\begin{split}&V^{1}_{23}=V_{0}\sqrt{k_{1}k_{2}k_{3}}W^{1}_{23}\,,\quad V_{0}=\frac{3\sqrt{c_{s}}}{4\sqrt{2}}\,,\\ &W^{1}_{23}=\frac{1}{2\sqrt{\eta_{1}\eta_{2}\eta_{3}}}+\frac{\sqrt{\eta_{1}\eta_{2}\eta_{3}}}{6k_{1}k_{2}k_{3}}\left(\frac{k_{1}^{3}}{\eta_{1}}-\frac{k_{2}^{3}}{\eta_{2}}-\frac{k_{3}^{3}}{\eta_{3}}\right)\,,\end{split}$ (6)
where $\eta_{i}\equiv\eta(k_{i})=\sqrt{1+(k_{i}\xi)^{2}/2}$ for $i=1,2,3$. The Bogoliubov WKE (6) was first derived in Dyachenko et al. (1992) in a slightly different form with mistakes in the prefactor. Finally, by assuming isotropy and performing angle average, the WKE becomes
$\frac{\partial n_{k}}{\partial t}=\mathrm{St}_{k}=4\pi^{2}V_{0}^{2}\int\limits_{k_{1}\,,k_{2}\geq 0}\big{(}|W^{k}_{12}|^{2}\mathcal{T}^{k}_{12}\delta(\omega^{k}_{12})$ (7)
@@
where $\mathcal{T}^{k}_{12}=n_{k_{1}}n_{k_{2}}-n_{k}n_{k_{1}}-n_{k}n_{k_{2}}$. In Appendix A, a revised, rigorous derivation of the WKE, its asymptotic regimes, and the parametrization of the resonant manifold can be found.
In this Letter, we obtain new theoretical predictions for the 3D Bogoliubov WTT in both large- and small-scale limits. We test these predictions by performing numerical simulations of the GPE (2) and the isotropic WKE (7).
In the large-scale limit $k\xi\to 0$, we have $\eta_{i}\to 1$ and therefore $\omega_{k_{i}}=c_{s}k_{i}$ – this is the acoustic wave limit. Also, on the resonant manifold we have $W^{1}_{23}\to 1$, and thus $V^{1}_{23}\to V_{0}\sqrt{k_{1}k_{2}k_{3}}$, consistently with the result in Griffin et al. (2022). The collisional integral of the WKE (7) becomes as in the 3D acoustic WKE Zakharov and Sagdeev (1970),
@@
Reference Zakharov and Sagdeev (1970) obtained a Kolmogorov-type stationary solution for the energy spectrum of 3D acoustic waves: $E(k)=4\pi k^{2}\omega_{k}n_{k}=C_{1}\,c_{s}^{1/2}P_{0}^{1/2}k^{-3/2}$, the so-called Zakharov-Sagdeev (ZS) spectrum. Here, the energy flux is defined as $P(k)=-\int_{0}^{k}4\pi\,q^{2}\omega(q)\mathrm{St}(q)\mathrm{d}q$, which is constant due to the spectrum’s stationarity, $P=P_{0}=$ const. An approximation of the dimensionless constant $C_{1}$ was given in Zakharov and Sagdeev (1970), as well in a recent paper Kochurin and Kuznetsov (2024). We derive its exact analytical expression in Appendix B. The complete ZS spectrum reads:
$E(k)=C_{1}c_{s}^{1/2}P_{0}^{1/2}k^{-3/2}\,,C_{1}=1/\sqrt{3\pi(\pi+4\ln{2}-1)}\,.$ (9)
Note that the acoustic WKE (8) is derived as an asymptotic limit of (7). Unlike the 2D case where $St_{k}\propto\xi^{-1}$ Griffin et al. (2022), in 3D, $\xi$ cancels out after angle averaging Zakharov et al. (1992) leading to (8). However, in practice, a small amount of dispersion is still crucial in the system to regularize the wave dynamics and avoid singularities. Note that in the recent work Kochurin and Kuznetsov (2024), the authors numerically obtained the ZS spectra in 3D acoustic wave systems, and different values of constant were reported for the dispersive and non-dispersive cases. In Appendix A and B, we show that the WKE of non-dispersive acoustic waves is regular for the 3D isotropic case, but with a different prefactor. Consequently, the ZS spectrum comes with a different constant.
In the small-scale limit $k\xi\to\infty$, we have $\eta_{k}\to k\xi/\sqrt{2}$ and $V^{1}_{23}\to 2^{7/4}V_{0}\xi^{-3/2}/3$. The collisional integral now becomes
@@
where now $\omega_{k}=c\xi k^{2}/\sqrt{2}$. Note that this collision integral coincides with the one for the finite-$k$ spectral part $n_{k}^{\prime}$ obtained from the four-wave kinetic equation via substitution $n_{k}=N_{0}\delta({\bf k})+n_{k}^{\prime}$, $N_{0}=$ const; see e.g. Düring et al. (2009). This fact is natural because both imply weak condensate, and the four-wave kinetic equation additionally assumes that the condensate’s phase is random.
To find the Kolmogorov-type spectrum in the short-wave limit, we apply the standard procedure provided by the WTT and seek a stationary power-law solution of (10), $n_{k}=Ak^{-x},\;A=$ const. This leads to $\mathrm{St}_{k}=64V_{0}^{2}\pi^{2}A^{2}(9\xi^{4}c_{s})^{-1}k^{1-2x}I(x)$, with the dimensionless collision integral
@@
The energy flux is $P(k)=\frac{128\sqrt{2}}{9}\pi^{3}V_{0}^{2}\xi^{-3}A^{2}\frac{I(x)}{2x-6}k^{6-2x}$. Constant energy flux, $P(k)\equiv P_{0}=$ const, implies $02x-6=$ 0, i.e. $x=3$ and $I(3)=0$. Using L’Hôpital rule, $\lim_{x\to 3}I(x)/(x-3)=dI(x)/dx|_{x=3}$. In Appendix B, we evaluate this expression and find
$E(k)=C_{2}c_{s}^{1/2}\xi^{5/2}P_{0}^{1/2}k\,,C_{2}=2^{3/4}/\sqrt{\pi(\pi-4\ln{2})}\,.$ (12)
@@
This spectrum, as well as the constant $C_{1}$ in (9), are new theoretical predictions. Note that the constants $C_{1}$ and $C_{2}$ are finite because the respective integrals defining the energy flux are convergent – a property called locality of the wave interactions (see Appendix B). We will now test these predictions numerically via simulations of the GPE and the isotropic WKE.
We perform numerical simulations of the forced-dissipated GPE (2) using the standard massively-parallel pseudo-spectral code FROST Krstulovic (2020) with a fourth-order Exponential Runge-Kutta temporal scheme (see Zhu et al. (2022, 2023a, 2023b)). In addition, the condensate amplitude (at $k=0$) is kept constant during the evolution. We set the initial data with uniform condensate with $|\Psi(0)|^{2}=\rho_{0}$, the forcing then adds the wave disturbances, and we evolve the system until it reaches a steady state. We then perform averages over time. We use grids of $N_{p}^{3}$ collocation points. We set density $\rho_{0}=1$ and $c_{s}=1$, and $\xi=\theta\Delta x$ where the grid length $\Delta x=L/N_{p}$ and $L=2\pi$. The forcing term is supported on a narrow band ($1\leq k\leq 3$) and it obeys the Ornstein–Ulenbeck process $\mathrm{d}F_{\bf k}(t)=-\gamma\,\widehat{\Psi}_{\bf{k}}\mathrm{d}t+f_{\rm w}\mathrm{d}{W}_{\bf k}$ where $W_{\bf k}$ is the Wiener process. The parameters $\gamma$ and $f_{0}$ control the correlation time and the amplitude of the forcing respectively. Dissipation is of the form $D_{\bf k}=(k/k_{\rm R})^{2\alpha}$ acting at the small scales. Finally, the $k$-space energy fluxes $P(k)$ are computed directly using the GPE (2) (see SM of Griffin et al. (2022)). We verify that the simulated WT is in a weakly nonlinear regime by measuring spatio-temporal spectra of $\Psi({\bf x},t)$ and confirming that they are narrowly concentrated around the dispersion curve (1) (see Appendix C).
We simulate the isotropic WKE (7) and (8), using the code WavKinS Krstulovic and Labarre (2024). We also include forcing and dissipation terms, similarly to the GPE. We use a logarithmic grid of $N_{r}$ points spanning the spectral domain $[k_{\rm min},k_{\rm max}]$. The collision integral is a 1D integral obtained by making explicit use of the parametrization of the resonant manifold provided in Appendix A. It is computed using the trapezoidal rule on the log-grid and a linear interpolation scheme. Time advancement is done by the Runge-Kutta 2 method. Forcing is of the form $f_{k}=f_{0}k^{s}e^{-(k-k_{f})^{2}/\Delta k_{f}^{2}}$ and dissipation $-(k/k_{d})^{\beta}n_{k}$. Parameters are optimized to obtain a large inertial range, while resolving the spectrum well. See Appendix C for more details on our methods.
@@
Figure 1: Numerical results obtained by simulating WKE and GPE respectively. (a) Dimensionless energy spectra $E(k)/\sqrt{c_{s}P_{0}\xi^{3}}$vs. dimensionless wave number $k\xi$ with theoretical predictions (9) and (12) superimposed. (b) Corresponding energy fluxes normalized by their values measured in the inertial range.
@@
We perform two isotropic WKE simulations: one of the Bogoliubov WKE (7) and the other one of the purely acoustic WKE (8). We also perform two GPE simulations with $\theta=1.25$ and $\theta=15$ to focus on different ranges concerning the transitional wave number $k_{\xi}=1/\xi$. The results of these runs in terms of the energy spectra and the fluxes superimposed with the respective theoretical predictions (9) and (12) are presented in Fig 1. The energy fluxes shown in Fig 1(b) exhibit plateau ranges, with $P(k)=P_{0}=$ const, in which the spectra are stationary and both forcing and dissipation effects are negligible (the so-called inertial range). The measured values of $P_{0}$ are then used for normalizing the energy spectra in Fig 1(a). We can see that each of the respective inertial ranges, for both WKE and GPE simulation, are in excellent agreement with our theoretical predictions (both in the spectral slopes and the prefactor constants). The longest inertial range, spanning for about two decades, is realized in the acoustic WKE run with scaling spectrum (9) confirmed. The Bogoliubov WKE run produced simultaneously two scaling ranges in the large-scale and the short-scale parts of the spectrum (about one decade wide each). The GPE runs with small and big values of $\xi$ produced one scaling each for $k\xi<1$ and $k\xi>1$. The scaling range in the GPE run with $\theta=1.25$ is the narrowest among all the simulations, but its spectrum agrees with the Bogoliubov WKE simulation. In particular, it exhibits a bump at the high-$k$ part of the spectrum close to the transitional wave number, $k\sim k_{\xi}$. A similar bump was previously observed in GPE simulations of Krstulovic and Brachet (2011), and it was called a ”dispersive bottleneck” since it is related to an increased wave dispersion in this region of scales.

### v4

@@ Abstract
Bogoliubov waves are fundamental excitations of Bose-Einstein Condensates (BECs). They emerge from a perturbed ground state and interact nonlinearly, triggering turbulent cascades. Here, we study turbulent BECs theoretically and numerically using the 3D Gross-Pitaevskii model and its associated wave-kinetic equations. We derive a new Kolmogorov-like stationary spectrum for short Bogoliubov waves and find a complete analytical expression for the spectrum in the long-wave acoustic regime. We then use our predictions to explain the BEC equation of state reported in [Dora et al. Nature 620,521 (2023)], and to suggest new experimental settings.
@@
One of the most interesting and important regimes in BEC turbulence is realized when weak waves propagate on a background of a strong coherent ground state $\Psi_{0}$, the condensate. These excitations are the Bogoliubov waves, whose frequency $\omega$ is given by the dispersion relation,
@@
where $k=|{\bf k}|$ is the modulus of the wave vector $\bf k$}, $c_{s}$ is the speed of sound, and $\xi$ is the healing length of the BEC at which dispersive effects become important. The wave field $\Psi(\bm{x},t)$ can be viewed as a ground condensate with superimposed perturbations that $\Psi(\bm{x,}t)=\Psi_{0}(\bm{x,t})+\delta\Psi(\bm{x,t})$. Bogoliubov waves emerge when the condensate amplitude $|\Psi_{0}|$ is significantly larger than the perturbations $|\delta\Psi|$, leading to quadratic nonlinearity and 3-wave interactions.
@@
Over the last few years, outstanding progress has been achieved in realizing turbulent BECs experimentally, and they have become an excellent platform for studying far-from-equilibrium physics. Turbulent states are driven by external forcing and dissipation, which results in an energy flux $P_{0}$ through scales (which equals the energy injection and dissipation rates). Such fluxes thus play a role equivalent to that of temperature in equilibrium thermal states. It is then natural to ask if there is a universal far-from-equilibrium equation of state (EoS) that relates energy flux and internal observables, such as energy or particle momentum distribution. The clean and well-controlled turbulent BEC experiments available today Dogra et al. (2023), together with recent numerical simulations of the Gross-Pitaevskii equation Martirosyan et al. (2024), have led to revealing the existence of such EoS. They have shown that the amplitude of the wave-action spectrum scales with the flux as $P_{0}^{1/3}$, followed by a scaling $P_{0}^{2/3}$-scaling for large flux. Whereas the $P_{0}^{1/3}$-scaling can be understood within the WTT invoking 4-wave interactions, the second scaling-law lacks explanation.
In the first part of this paper, we obtain new theoretical predictions for 3D BEC turbulence dominated by interacting Bogoliubov waves, which are then confirmed by our numerical simulations. Among the most significant outcomes of our work are the following stationary energy spectra of Kolmogorov-Zakharov (KZ) type, which include the universal constants, for the acoustic and short wave limits, respectively:
$E(k)$ $=C_{1}c_{s}^{1/2}P_{0}^{1/2}k^{-3/2},$ $\text{for }k\xi\ll 1,$ (2)
$E(k)$ $=C_{2}c_{s}^{1/2}\xi^{5/2}P_{0}^{1/2}k,$ $\text{for }k\xi\gg 1,$ (3)
where the universal constants are derived as $C_{1}=1/{\sqrt{3\pi(\pi+4\ln{2}-1)}}$ and $C_{2}={2^{3/4}}/{\sqrt{\pi(\pi-4\ln{2})}}$. As an application of our theory, in the second part of the paper, we provide a plausible theoretical explanation of the second scaling law of the EoS observed experimentally in Dogra et al. (2023). This experiment used a cylinder-shaped box trap that was shaken to excite turbulence and which had a ”finite height” thereby providing a dissipation (escape) of high-momentum particles. We re-analyze the experimental data of Dogra et al. (2023), and show their agreement with our prediction of scaling $P_{0}^{1/2}$ arising from interacting short-wave Bogoliubov modes. Our findings are important for BECs in a broad context, and as they are universal and expected to be generically realizable in turbulent systems, their impact thus goes beyond the explanation of a particular experiment.
Wave Turbulence Predictions for Bogoliubov waves
Stationary spectra of Komolgrov-Zakharov type
@@
The master model for describing BECs is the Gross-Pitaevskii equation (GPE) for the complex scalar wave function $\Psi({\bf x},t)$, where $\bf x$} is the position in 3D physical space and $t$ is time. Defining $\Psi$ such that $|\Psi|^{2}$ is the mass density, the GPE, written in terms of $\xi$, $c_{s}$ and the unperturbed (uniform) condensate mass density $\rho_{0}$, reads
$i\frac{\partial\Psi}{\partial t}=\frac{c_{s}}{\sqrt{2}\xi}\Big[-\xi^{2}\nabla^{2}+\frac{|\Psi|^{2}}{\rho_{0}}-1\Big]\Psi+V(\bm{x},t)\Psi+i\,\mathcal{F}-i\,\mathcal{D}\,.$ (4)
We have also included an external potential trap $V(\bm{x})$, a large-scale forcing $\mathcal{F}$ and a hyper-viscous dissipation term $\mathcal{D}$ acting at small scales. In BEC experiments, dissipation is achieved synthetically by allowing high energy particles to escape from the trap thereby removing energy from the system. In our simulations we mimic this effect by the hyper-viscous term which effectively acts on the high-$k$ particles only. Moreover, in experiments, energy injection is provided by shaking the confining trap. First, in our idealized simulations, we set $V\equiv 0$, but we include an external forcing in order to obtain an non-equilibrium steady state. Then, we verify our predictions in a trapped BEC, with an oscillating trap (and $\mathcal{F}=0)$. The healing length and the speed of sound depend both on the physical properties of the BEC and on $\rho_{0}$.
@@
The GPE can be mapped to an effective compressible irrotational fluid flow via the so-called Madelung transformation, $\Psi({\bf x},t)=\sqrt{\rho({\bf x},t)}\exp[i\phi({\bf x},t)/\sqrt{2}c_{s}\xi]$ with $\rho({\bf x},t)$ and $\phi({\bf x},t)$ being the fluid mass density and the velocity potential, respectively. The Bogoliubov dispersion relation (1) is obtained by linearizing the GPE, (4), in the absence of the potential, forcing, and dissipation terms, considering perturbations over the ground state, $\delta\Psi=\Psi-\sqrt{\rho_{0}}$, and diagonalizing the linearized equations in Fourier space.
In our theory, we consider scales much smaller than the size of the potential trap, which allows us to consider the large-box limit and $V(\bm{x})\equiv 0$. Then, for weakly nonlinear waves, the leading order interaction occurs between the modes satisfying 3-wave resonant conditions,
@@
In this limit, the WTT provides a wave-kinetic equation (WKE) describing the evolution of the wave-action spectrum $n^{\text{B}}_{{\bf k}}=n^{\text{B}}({{\bf k}},t)$ which is the Bogoliubov quasi-particle momentum distribution defined in (17) in Methods. The general WKE for weakly nonlinear waves driven by 3-wave resonant interactions is Zakharov et al. (1992); Nazarenko (2011)
$\frac{\partial n^{\text{B}}_{{\bf k}}}{\partial t}=\mathrm{St}_{{\bf k}},$ (6)
@@
[…]
where $V^{k}_{{12}}\equiv V({k},{k_{1}},{k_{2}})$ is the 3-wave interaction amplitude. For Bogoliubov waves
[…]
where $\eta_{i}\equiv\eta(k_{i})=\sqrt{1+(k_{i}\xi)^{2}/2}$ for $i=1,2,3$. The Bogoliubov WKE (8) was first derived in Dyachenko et al. (1992) in a slightly different form with mistakes in the prefactor. Finally, by assuming isotropy and performing an angle average, the WKE becomes
$\frac{\partial n^{\text{B}}_{k}}{\partial t}=\mathrm{St}_{k}=4\pi^{2}V_{0}^{2}\int\limits_{k_{1}\,,k_{2}\geq 0}\big(|W^{k}_{12}|^{2}\mathcal{T}^{k}_{12}\delta(\omega^{k}_{12})$ (9)
@@
where $\mathcal{T}^{k}_{12}=n^{\text{B}}_{k_{1}}n^{\text{B}}_{k_{2}}-n^{\text{B}}_{k}n^{\text{B}}_{k_{1}}-n^{\text{B}}_{k}n^{\text{B}}_{k_{2}}$ (Methods). Note that only resonant waves interact.
Acoustic limit
In the large-scale limit $k_{i}\xi\to 0,\;$ we have $\eta_{i}\to 1$ and therefore $\omega_{k_{i}}=c_{s}k_{i}$ – this is the acoustic-wave limit. Also, on the resonant manifold (5), we have $W^{1}_{23}\to 1$, and thus $V^{1}_{23}\to V_{0}\sqrt{k_{1}k_{2}k_{3}}$, consistently with the result in Griffin et al. (2022). The collisional integral of the WKE (9) becomes as in the 3D acoustic WKE Zakharov and Sagdeev (1970),
@@
In Zakharov and Sagdeev (1970), Zakharov and Sagdeev obtained a Kolmogorov-type stationary solution for the energy spectrum of 3D acoustic waves: $E(k)=4\pi k^{2}\omega_{k}n^{\text{B}}_{k}=C_{1}\,c_{s}^{1/2}P_{0}^{1/2}k^{-3/2}$, the so-called Zakharov-Sagdeev (ZS) spectrum. Here, the energy flux $P_{0}$ is constant because of the stationarity of the spectrum. An approximation of the dimensionless constant $C_{1}$ was given in Zakharov and Sagdeev (1970), as well in a recent paper Kochurin and Kuznetsov (2024). We derive its exact analytical expression, and obtain the complete ZS spectrum, (2) (Methods).
Note that the acoustic WKE (10), is derived as an asymptotic limit of (9). Unlike the 2D case where $St_{k}\propto\xi^{-1}$ Griffin et al. (2022), in 3D $\xi$ cancels out after angle averaging Zakharov et al. (1992) leading to (10). However, for its validity, a small amount of dispersion, naturally present in BECs, is still crucial in the system to regularize the wave dynamics and avoid singularities.
Short-wave limit
In the small-scale limit $k\xi\to\infty$, we have $\eta_{k}\to k\xi/\sqrt{2}$ and $V^{1}_{23}\to 2^{7/4}V_{0}\xi^{-3/2}/3$. The collisional integral becomes
@@
where now $\omega_{k}=c\xi k^{2}/\sqrt{2}$. Note that this collision integral coincides with the one for the finite-$k$ spectral part $n^{\text{B}}_{k}$$^{\prime}$ obtained from the 4-wave kinetic equation via substitution $n^{\text{B}}_{k}=N_{0}\delta({\bf k})+{n^{\text{B}}_{k}}^{\prime}$, $N_{0}=$ const; see e.g. Düring et al. (2009).
To find the Kolmogorov-type spectrum in the short-wave limit, we apply the standard procedure provided by the WTT and seek a stationary power-law solution of (11), $n^{\text{B}}_{k}=Ak^{-x},\;A=$ const. This leads to $\mathrm{St}_{k}=64V_{0}^{2}\pi^{2}A^{2}(9\xi^{4}c_{s})^{-1}k^{1-2x}I(x)$, with the dimensionless collision integral
@@
The standard definition of the energy flux is given by $P(k)=-\int_{0}^{k}4\pi\,q^{2}\omega(q)\mathrm{St}(q)\mathrm{d}q=\frac{128\sqrt{2}}{9}\pi^{3}V_{0}^{2}\xi^{-3}A^{2}\frac{I(x)}{2x-6}k^{6-2x}$. Constant energy flux, $P(k)\equiv P_{0}=$ const, implies 2x6=02x-6=0, i.e. $x=3$ and $I(3)=0$. Using the L’Hôpital rule, $\lim_{x\to 3}I(x)/(x-3)=dI(x)/dx|_{x=3}$. We evaluate this expression and find the steady spectrum, (3) (Methods). This spectrum, as well as the constant $C_{1}$ in (2), are new theoretical predictions. Note that the constants $C_{1}$ and $C_{2}$ are finite because the respective integrals defining the energy flux are convergent, a property called locality of the wave interactions (Methods). We will now test these predictions numerically via simulations of the GPE and the isotropic WKE.
Numerical Validation
@@
Figure 1: Numerical results obtained by simulating WKE and GPE. (a) Dimensionless energy spectra $E(k)/\sqrt{c_{s}P_{0}\xi^{3}}$ vs. dimensionless wave number $k\xi$ with theoretical predictions (2) and (3) superimposed. Run 1 is performed using the WKE (9) and contains both the acoustic (long-wave) and short-wave regimes, whereas run 2 uses the $3D$ acoustic WKE (10). The GPE runs 3 and 4 are obtained for homogeneous cases with stochastic forcing, and runs 5 and 6 are obtained with a shaken potential trap. All the GPE runs are optimized such that the inertial range extends through the scales of interest. (b) Corresponding energy fluxes of run 1-4 normalized by their values measured in the inertial range. (c) Snapshot of density for GPE run 5 to verify the prediction in the acoustic limit. (d) Snapshot of density for GPE run 5 to verify the prediction in the short-wave limit.
First, we numerically simulate the forced-dissipated GPE, (4), setting $\rho_{0}=1$ and $c_{s}=1$, in a triply periodic cube with the side length $L=2\pi$, and $V(\bm{x})\equiv 0$. To focus on different ranges with respect to the transitional wavenumber $k_{\xi}=1/\xi$, two high-resolution GPE simulations with $\xi=1.25\Delta x$ and $\xi=15\Delta x$ ($\Delta x$ is the length of the grid spacing) are performed. Such choices allow us to study the acoustic and the short-wave limits separately. We set the initial data with uniform condensate with $|\Psi(0)|^{2}=\rho_{0}$, a stochastic forcing then adds the wave disturbances, and we evolve the system until it reaches a steady state, balancing between the large-scale forcing and the small-scale dissipation (Methods). We then perform averages over time. Also, we perform two isotropic WKE simulations: one of the Bogoliubov WKE, (9), and the other of the purely acoustic WKE, (10), with extra forcing and dissipation terms similar to the ones used in the GPE simulations (Methods).
@@
The results of these runs in terms of the energy spectra and the fluxes superimposed with the respective theoretical predictions, (2) and (3), are presented in Fig 1. The energy fluxes shown in Fig 1(b) exhibit plateau ranges, with $P(k)=P_{0}=$ const, in which the spectra are stationary and both forcing and dissipation effects are negligible (the so-called inertial ranges). The measured values of $P_{0}$ are then used for normalizing the energy spectra in Fig 1(a). We can see that each of the respective inertial ranges, for both WKE and GPE simulations, are in excellent agreement with our theoretical predictions (both for the spectral slopes and the prefactor constants). The longest inertial range, spanning for about two decades, is realized in the acoustic WKE run (run 2) with scaling spectrum (2) confirmed. The Bogoliubov WKE run (run 1) produced two scaling ranges in the large-scale and the short-scale parts of the spectrum simultaneously (about a decade-wide each). The GPE runs with small and big values of $\xi$ produced one scaling each for $k\xi<1$ (run 3) and $k\xi>1$ (run 4) . The scaling range in the GPE run 4 is the narrowest among all the simulations, but its spectrum agrees with the Bogoliubov WKE simulation. In particular, it exhibits a bump in the high-$k$ part of the spectrum close to the transitional wave number, $k\sim k_{\xi}$. A similar bump was previously observed in GPE simulations of Krstulovic and Brachet (2011), and it was called a ”dispersive bottleneck” since it is related to an increased wave dispersion in this region of scales. Note that our predictions are well confirmed by the GP simulations without any adjustable parameters.
In BEC turbulence experiments, the cloud is always confined by an external potential trap. To be more realistic, we perform two further GPE simulations in a box-shaped potential trap of size $L_{\rm trap}=1.5\pi$, with $\xi=1.25\Delta x$ (run 5) and $\xi=10\Delta x$ (run 6), namely $L_{\rm trap}=230.4\,\xi$ and $L_{trap}=28.8\,\xi$, respectively. We prepare ground states so that the initial density inside the trap is nearly constant with $|\Psi(0)|^{2}=\rho_{0}$. After that, we shake the trap in both the $x$ and $y$ directions to introduce energy into the system until the small-scale energy dissipation rate becomes equal to the injection rate. We adjust the shaking protocol and the dissipation parameters to keep the system in the Bogoliubov regime ensuring that at least 70% of the particles are staying at $k\leq k_{trap}=2\pi/L_{\rm trap}$. We measure the energy flux $P_{0}$ by calculating the dissipation rate of energy on small scales, in the same spirit as in Dogra et al. (2023) (Methods). Fig 1(c) and (d) show the density snapshots for runs 5 and 6, respectively. Although the presence of traps reduces the width of the inertial range, Fig. 1(a) shows that the scalings of the stationary spectrum predicted by WWT survived with the potential trap for both the acoustic and short-wave limits. The interaction with the potential trap increases the spectrum constants: namely, we observe that the spectra obtained with the potential trap are both somewhat greater than those corresponding to GPE simulations without the potential trap, but they still agree well with the predicted $k$-scaling and the order of magnitude of their amplitudes
Interpreting the experimental EoS
EoS predicted by the 4-wave and the 3-wave WT